# Thermodynamic and energetic constraints on out-of-equilibrium tunneling rates

### v1

@@ III.2 Energetic constraint
Similarly to the thermodynamic constraint, Eq. (20) directly translates into an energetic constraint on the integral over the out-of-equilibrium noise in Eq. (8b) by taking the sum of the energetic constraints for both tunneling directions. However, this integral makes the energetic constraint less relevant from an experimental point of view because it requires knowledge of both internal energy and tunneling rates at all intermediate temperatures of subsystem L.
Note that the derivations of both thermodynamic and energetic constraints do in principle not require the temperatures of the two subsystems to be equal neither at the beginning nor at end of any of the strokes (see Appendix D and F). Here, however, we choose one of the settings where the two subsystems have equal temperature, in order to be able to compare to an easily accessible, experimentally relevant reference situation when establishing the out-of-equilibrium tunneling-rate constraints.
III.3 Thermodynamic limit
Here, we comment on the relevance of the discovered constraints (16) and (20) for subsystems that approach the thermodynamic limit. Notably, both the thermodynamic and the energetic constraints combine transport quantities, i.e., the tunneling rates, with extensive properties of subsystem L, namely its internal energy and its nonequilibrium free energy. However, the extensive quantities only appear on the left-hand side of Eqs. (16, 20), and not on the right-hand side. This feature implies that the left- and the right-hand sides behave very differently depending on the size of the system. More concretely, if the extensive quantities in the thermodynamic limit scale as
$U_{\text{L}}\to\lambda U_{\text{L}},\qquad\Delta F_{\text{L}}^{\text{(c,h)}}\to\lambda\Delta F_{\text{L}}^{\text{(c,h)}}\ ,$ (21)
where $\lambda$ is the scale parameter, and the tunneling rates scale with an arbitrary scaling function $f(\lambda)$, i.e. $\Gamma_{\rightleftarrows}\to f(\lambda)\Gamma_{\rightleftarrows}$, both the thermodynamic constraint of Eq. (16) and the energetic constraint of Eq. (20) become trivial. Indeed, for $\lambda\to\infty$, the right-hand side containing the rate response $\mathcal{W}^{(\text{Resp})}_{\rightleftarrows}$ becomes negligible, and Eqs. (16, 20) then reduce to expressions that state the positivity of the thermodynamic and energetic costs, respectively, on the left-hand side. For instance, we can imagine replacing subsystem L with $\lambda$ copies of it, but take only one of the copies to be connected to subsystem R through the tunneling Hamiltonian $\hat{V}(t)$. Then, the internal energy and the entropy of the $\lambda$ copies are given as $U_{\lambda\text{L}}(\beta)=\lambda U_{\text{L}}(\beta)$ and $S_{\lambda\text{L}}(\beta)=\lambda S_{\text{L}}(\beta)$, respectively. This linear scaling implies that also the nonequilibrium free energy scales linearly with $\lambda$, i.e., $\Delta F_{\lambda\text{L}}^{\text{(c,h)}}=\lambda\Delta F_{\text{L}}^{\text{(c,h)}}$. By stark contrast, both the temperatures $T_{\rm L}$, $T_{\rm R}$ and the tunneling rates remain unchanged because the former are intensive properties and the latter are transport properties that involve only one of the copies. Therefore, in the thermodynamic limit, the left-hand sides of both thermodynamic and energetic constraints of Eqs. (16, 20) scale with $\lambda$ whereas the right-hand sides do not.
However, the fact that the right hand sides of these equations can be neglected in the thermodynamic limit—independently on how it is taken—leads to a trivial statement. The thermodynamic and energetic costs are positive by construction: The tunneling rates are positive $\Gamma_{\rightleftarrows}\geq 0$, and so are the coefficients of performance $\eta^{\text{(c,h)}}\geq 0$, the nonequilibrium free energies $-\Delta F_{\text{L}}^{\text{(c,h)}}\geq 0$, as well as the energy variation $dU_{\text{L}}(x)$ in the integral in Eq. (20).
Thus, for both the thermodynamic and the energetic constraint to be insightful, subsystem L (namely the subsystem which we consider at two different temperatures in order to compare tunneling rates) must be a small quantum system, namely it must not satisfy the thermodynamic limit.
IV Examples
Figure 3: Illustration of the two setups considered in Sec. IV. (a) A two-level atom with frequency $\omega_{\text{a}}$ is weakly coupled, $g\ll\hbar\omega_{\text{a,c}}$, to a cavity with characteristic frequency $\omega_{\text{c}}$ and Kerr non-linearity $u_{c}$. (b) Two interacting fermionic tight-binding rings with different sizes are weakly coupled at a single (“0”) site.
To illustrate the constraints derived in Sec. III, we consider two different physical settings, represented by the systems depicted in Fig. 3: An atom coupled to a nonlinear cavity [panel (a)], and two fermionic chains interchanging particles at a single site [panel (b)]. Such examples are chosen not only because of illustration purposes, but also because they are of experimental relevance to test our predictions with state-of-the art setups. These two examples include optical or mechanical cavities coupled together [61] or to (artificial) atoms [62, 63, 64, 65, 66, 67], and tunneling bridges across molecules [68, 38, 40] or magnetic impurities [69, 70, 71].
@@ IV Examples
To clarify the constraints of Eqs. (16, 20), we notice that both the thermodynamic, $\mathcal{W}_{\rightleftarrows}^{\text{(Thermo)}}$, and energetic cost, $\mathcal{W}_{\rightleftarrows}^{\text{(Energy)}}$, as well as the rate response $\mathcal{W}_{\rightleftarrows}^{\text{(Resp)}}$ entering the constraints can be recast as sums over resonances as
@@
$\mathcal{W}^{(i)}_{\pm}=\sum_{nmlk}w^{(i)}_{\pm,nmlk}\delta(\epsilon_{mn}^{\text{(L)}}+\epsilon_{kl}^{\text{(R)}}-\hbar\omega),$ (22)
@@
and thus sum over all the resonances at the same frequency $\omega$. Since both the thermodynamic and energetic constraints hold separately at each resonance, Eqs. (16, 20) also hold for each amplitude $\mathcal{C}_{\rightleftarrows}^{(i)}(\omega)$, i.e.
$\begin{split}\mathcal{C}_{\rightleftarrows}^{\text{(Thermo)}}(\omega)&\geq\mathcal{C}_{\rightleftarrows}^{\text{(Resp)}}(\omega),\\ \mathcal{C}_{\rightleftarrows}^{\text{(Energy)}}(\omega)&\geq\mathcal{C}_{\rightleftarrows}^{\text{(Resp)}}(\omega),\\ \end{split}$ (24)
which establishes natural quantities to compute in our analysis of the systems depicted in Fig. 3.
IV.1 Atom coupled to nonlinear cavity
Figure 4: Thermodynamic cost (dashed lines), energetic cost (dotted lines), and rate response (solid lines) for an atom coupled to a cavity as functions of the inverse temperature difference $\beta_{\text{a}}-\beta_{\text{c}}$. The atom and cavity frequencies are taken in resonance, $\omega_{\text{a}}=\omega_{\text{c}}$, and the different curve colors denote different Kerr nonlinearities, $u_{\text{c}}/\hbar\omega_{\text{a}}=0,1,10$, in blue, green, and red, respectively. In panels (a, c) the cavity is kept at the cold temperature ($\beta_{\text{c}}>\beta_{\text{a}}$), whereas in panels (b, d) it is the atom that has the colder temperature ($\beta_{\text{a}}>\beta_{\text{c}}$).
@@ IV.1 Atom coupled to nonlinear cavity
Here, one of the subsystems is the atom, described by the Pauli matrix $\hat{\sigma}_{z}$ and has the atom frequency $\omega_{\text{a}}$. The cavity is the other subsystem, and is described by the number operator $\hat{n}=\hat{a}^{\dagger}\hat{a}$, with $[\hat{a},\hat{a}^{\dagger}]=1$, and the cavity frequency $\omega_{\text{c}}$. We also include a Kerr nonlinearity, parametrized by $u_{\text{c}}$, which plays the role of effective interactions between cavity photons in this system. The atom and the cavity exchange photon quanta through the weak tunneling term $g(\hat{a}\hat{\sigma}_{+}+\hat{a}^{\dagger}\hat{\sigma}_{-})$ with $g\ll\hbar\omega_{\text{a,c}}$ and where $\hat{\sigma}_{+}$ and $\hat{\sigma}_{-}$ are the raising and lowering operators of the atom states. Note that we choose $\omega=0$ since any external driving frequency can be incorporated in the laser detuning—replacing $\omega_{c}$—in the rotating frame [42]. For $u_{\text{c}}=0$, the Hamiltonian (25) reduces to the Jaynes-Cummings Hamiltonian [73].
@@
To relate to the theoretical framework of Sec. II, we now identify the tunneling operator $\hat{A}=g\hat{a}\hat{\sigma}_{+}$, which we use to calculate the zero-frequency amplitudes $\mathcal{C}_{\rightleftarrows}^{(i)}(0)$ in Eq. (23). Note that a possible observable $\hat{Q}$ could here be the atomic occupation $\hat{Q}=\frac{\hat{\sigma}_{z}+1}{2}$, satisfying $[\hat{Q},\hat{A}]=\hat{A}$. The result for $\mathcal{C}_{\rightleftarrows}^{(i)}(0)$ is shown in Fig. 4, where we plot $\mathcal{C}_{\rightleftarrows}^{(i)}(0)$ for different values of the non-linearity $u_{\text{c}}$. Since we are free to choose the L subsystem, in panels (a, c) the atom is the subsystem considered at two different temperatures, whereas in panels (b, d) the cavity is. Here we see that both thermodynamic and energetic constraints are trivially saturated at equal temperature. However, for sizable temperature biases, considering the atom or the cavity at different temperatures affects the constraints. In particular, in panels (b, d) we see that the cavity non-linearity allows to approach both thermodynamic and energetic constraints also at large temperature biases for both rates $\Gamma_{\rightleftarrows}$. The non-linearity affects the constraints (24) in two ways: On the one hand, finite values of $u_{\text{c}}$ break the degeneracy of the atom-cavity transitions, making the atom couple only to two consecutive cavity states. On the other hand, a large $u_{\text{c}}$ increases the energy spacing of the cavity. This feature reduces the number of states that have non-negligible occupation at finite cavity temperature $\beta_{\text{c}}\neq 0$. These two aspects effectively reduce the dimension of the cavity, moving it further away from the thermodynamic limit.
@@
By contrast, when it is instead the atom that is taken at two different temperatures, illustrated in panels (a, c), the rate $\Gamma_{\rightarrow}$ has a nontrivial constraint for all temperature biases, while the rate $\Gamma_{\leftarrow}$ does not. This stems from the transitions allowed by the tunneling operator $\hat{A}$ in the reverse process, which here consists of only the process in which the atom emits a photon into the cavity, combined with the energy spectrum of such allowed transitions. Specifically, if all allowed transitions start in states with energies that are more than a standard deviation away from the internal energy of the subsystem, here the atom, then the rate response is negative $\mathcal{W}^{\text{(Resp)}}\leq 0$, and the constraints become trivial, see Appendix G.
In Fig. 4, we also see that the thermodynamic cost and the energetic cost are very similar, and that there is no hierarchy between them, namely it depends on the specific parameters whether the thermodynamic or the energetic cost is larger. This can be seen in panel (b), where, in the absence of Kerr non-linearity the thermodynamic cost is larger than the energetic one, but in the presence of such a non-linearity the opposite holds true.
IV.2 Coupled fermionic rings
@@
Figure 5: Thermodynamic cost (dashed lines), energetic cost (dotted lines), and rate response (solid lines) for two (U and D) coupled fermionic rings with sizes $L_{\text{D}}=3$ and $L_{\text{U}}=10$. The costs are plotted vs the inverse temperature difference $\beta_{\text{D}}-\beta_{\text{U}}$. The ring hopping parameters, $t_{\text{D}}$ and $t_{\text{U}}$, and the driving frequency are fixed as $t_{\text{U}}=\frac{4}{5}t_{\text{D}}=2\hbar\omega$. The upper ring charging energy $u_{\text{U}}=0$ is also fixed, while the down ring charging energy is taken as $u_{\text{D}}/2t_{\text{D}}=0,1,10$ for curves in blue, green, and red, respectively. In panels (a, c) the upper, larger ring is kept at the cold temperature ($\beta_{\text{U}}>\beta_{\text{D}}$), whereas in panels (b, d) it instead is the down, smaller ring that is colder ($\beta_{\text{L}}>\beta_{\text{U}}$).
@@
Figure 6: Thermodynamic cost (dashed lines), energetic cost (dotted lines), and rate response (solid lines) for the two coupled fermionic rings with charging energies $u_{\text{D}}/2t_{\text{D}}=10,u_{\text{U}}=0$, plotted vs the inverse temperature difference $\beta_{\text{U}}-\beta_{\text{U}}$. The down ring has fixed size $L_{\text{L}}=3$, and the upper ring has sizes $L_{\text{U}}=4,7,10$ for the curves in blue, green, and red, respectively. All other parameters are the same as in Fig. 5. In panels (a, c), the upper, larger ring is kept at the cold temperature ($\beta_{\text{U}}>\beta_{\text{D}}$), whereas in panels (b, d) the down, smaller ring is kept fixed at the cold temperature ($\beta_{\text{D}}>\beta_{\text{U}}$).
@@
with $g\ll t$. Here, we identify the tunneling operator $\hat{A}=g\hat{c}^{\dagger}_{\text{D},0}\hat{c}_{\text{U},0}$, which transfers a fermion from the up ring (U) to the down ring (D). Then, using this tunneling operator $\hat{A}$, we calculate the amplitudes $\mathcal{C}_{\rightleftarrows}^{(i)}(\omega)$ of Eq. (23). Note that a possible observable $\hat{Q}$ could here be the number of fermions in one ring, e.g. $\hat{Q}=\hat{N}_{\text{D}}$, which satisfies $[\hat{Q},\hat{A}]=\hat{A}$. The result for $\mathcal{C}_{\rightleftarrows}^{(i)}(\omega)$ is shown in Fig. 5, where we plot $\mathcal{C}_{\rightleftarrows}^{(i)}(\omega)$ at different values of the charging energy $u_{\text{D}}$ and in Fig. 6 for different sizes $L_{\text{U}}$. Note that the frequency $\omega$ of the tunneling Hamiltonian can emerge from a potential bias between the rings after a gauge transformation [28, 74]. Similarly to the atom-cavity system considered in Sec. IV.1, the charging energy $u_{\text{D}}$ produces nontrivial constraints for both $\Gamma_{\rightleftarrows}$, as shown in panels (a, c) of Fig. 5, when the down ring is the one that is considered at two different temperatures when comparing out-of-equilibrium rates. By contrast, for vanishing charging energy $u_{\text{U}}=0$, only the rate $\Gamma_{\leftarrow}$ is non-trivially constrained, see panels (b, d), for the case where it is the upper ring that is taken at two different temperatures.
@@
Furthermore, as discussed in Sec. III.3, increasing the size of the ring that is considered at two different temperatures, when comparing out-of-equilibrium rates, weakens the constraints, as is seen in panels (b, d) of Fig. 6. In this example, the tunneling rates scale inversely with the ring size, i.e. with a scaling function $f$ on the form $f(L_{\text{U}})\sim 1/L_{\text{U}}$ [see below Eq. (21)], whereas the extensive quantities entering the thermodynamic and energetic costs scale linearly with $L_{\text{U}}$. Thus, while the costs are essentially unaffected upon increasing $L_{\text{U}}$, the rate responses instead decrease. By contrast, if the ring that is considered at two different temperatures has a fixed size, both costs and rate responses scale as $1/L_{\text{U}}$ because of the scaling of the tunneling rates alone, see panels (a, c). This feature implies that, as long as the size of the subsystem considered at two different temperatures remains small, there are nontrivial constraints on the tunneling rates, irrespective of the size of the subsystem with fixed temperature.
V Conclusions

### v2

@@ III.2 Constraints in the tunneling regime
We emphasize that Eq. (23a) directly translates into a constraint on the out-of-equilibrium noise, which is given by the sum of the tunneling rates via Eq. (22b). This also means, from a more practical, experimental viewpoint, that one possibility to test the constraint (23a) is by measuring the noise and tunneling current in two configurations: (i) When both subsystems have the same cold temperature, and (ii) in the desired out-of-equilibrium condition. One can then access the tunneling rates as $2q^{2}\Gamma_{\rightleftarrows}=\mathcal{S}\pm qI$. All the other quantities entering the inequality can be calculated once the state of the subsystem considered at the different temperatures is known  [^2] Note that it is not necessary to know the Hamiltonian of the subsystem kept at fixed temperature, which may be arbitrarily complicated, to test these constraints.. Similarly to the thermodynamic constraint, Eq. (23b) directly translates into an energetic constraint on the integral over the out-of-equilibrium noise in Eq. (22b) by taking the sum of the energetic constraints for both tunneling directions. However, this integral makes the energetic constraint less relevant from an experimental point of view because it requires knowledge of both internal energy and tunneling rates at all intermediate temperatures of subsystem L.
From the thermodynamic constraint (23a), we can also derive a direct lower bound on the out-of-equilibrium tunneling rates $\Gamma_{\rightleftarrows}(\beta_{\mathrm{h}},\beta_{\mathrm{c}})$ in terms of the equilibrium rates (see the derivation in Appendix E). These constraints furthermore contain the rate response, (integrals of) thermodynamic quantities, and they read
[…]
where we defined
$\begin{split}f(x)&\equiv\partial_{\text{L}}\Gamma_{\rightleftarrows}(\beta_{\mathrm{c}},\beta_{\mathrm{c}})-\Delta F_{\text{L}}^{\text{(c)}}(x)\eta^{\text{(h)}}(x)\Gamma_{\rightleftarrows}(\beta_{\mathrm{c}},\beta_{\mathrm{c}}),\\ g(x)&\equiv-\Delta F_{\text{L}}^{\text{(h)}}(x)\eta^{\text{(c)}}(x).\end{split}$ (26)
While the bound (25) has a more complex shape, containing integrals over thermodynamic functions, it has the important advantage that it does not depend on the out-of-equilibrium responses $\partial_{\text{L}}\Gamma_{\rightleftarrows}(\beta_{\mathrm{h}},\beta_{\mathrm{c}})$. Instead, only the more easily accessible equilibrium response function, $\partial_{\text{L}}\Gamma_{\rightleftarrows}(\beta_{\mathrm{c}},\beta_{\mathrm{c}})$, enters Eq. (25). We emphasize that Eq. (25) directly provides a lower bound for the out-of-equilibrium noise. In particular, compared to [25], it provides a nontrivial constraint indicating how much super-Poissonian the noise is.
@@
A further insightful way of writing the constraints of Eqs. (23a, 23b), is by highlighting their contributions from sums over resonances. Indeed, we notice that both the thermodynamic, $\mathcal{W}_{\rightleftarrows}^{\text{(Thermo)}}$, and energetic cost, $\mathcal{W}_{\rightleftarrows}^{\text{(Energy)}}$, as well as the rate response $\mathcal{W}_{\rightleftarrows}^{\text{(Resp)}}$ can be recast as
@@
$\mathcal{W}^{(i)}_{\leftrightarrows}=\sum_{nmlk}w^{(i)}_{\leftrightarrows,nmlk}\delta(\epsilon_{mn}^{\text{(L)}}+\epsilon_{kl}^{\text{(R)}}-\hbar\omega),$ (27)
@@
and thus sum over all the resonances at the same frequency $\omega$. Since both the thermodynamic and energetic constraints hold separately at each resonance, Eqs. (23a, 23b) also hold for each amplitude $\mathcal{C}_{\rightleftarrows}^{(i)}(\omega)$, i.e.
$\begin{split}\mathcal{C}_{\rightleftarrows}^{\text{(Thermo)}}(\omega)&\geq\mathcal{C}_{\rightleftarrows}^{\text{(Resp)}}(\omega),\\ \mathcal{C}_{\rightleftarrows}^{\text{(Energy)}}(\omega)&\geq\mathcal{C}_{\rightleftarrows}^{\text{(Resp)}}(\omega).\\ \end{split}$ (29)
This establishes natural quantities, which we will compute in the following analysis of the example systems.
Figure 4: Illustration of the two setups considered in Sec. III. (a) A two-level atom with frequency $\omega_{\text{A}}$ is weakly coupled, $g\ll\hbar\omega_{\text{A,C}}$, to a cavity with characteristic frequency $\omega_{\text{C}}$ and Kerr non-linearity $u_{\mathrm{C}}$, see Sec. III.3. (b) Two interacting fermionic tight-binding rings with different sizes are weakly coupled at a single (“0”) site, see Sec. III.4.
Concretely, to illustrate the constraints (23a) and (23b), we consider two different physical settings, represented by the systems depicted in Fig. 4: An atom coupled to a nonlinear cavity [panel (a)], see Sec. III.3, and two fermionic chains interchanging particles at a single site [panel (b)], see Sec. III.4. Such examples are chosen not only because of illustration purposes, but also because they are of experimental relevance to test our predictions with state-of-the art setups. These two examples include optical or mechanical cavities coupled together [46] or to (artificial) atoms [47, 48, 49, 50, 51, 52], and tunneling bridges across molecules [53, 37, 39] or magnetic impurities [54, 55, 56].
III.3 Atom coupled to nonlinear cavity
Figure 5: Thermodynamic cost (dashed lines), energetic cost (dotted lines), and rate response (solid lines) for an atom coupled to a cavity as functions of the inverse temperature difference $\beta_{\text{A}}-\beta_{\text{C}}$. The atom and cavity frequencies are taken in resonance, $\omega_{\text{A}}=\omega_{\text{C}}$, and the different curve colors denote different Kerr nonlinearities, $u_{\text{C}}/\hbar\omega_{\text{A}}=0,10$, in blue and red, respectively. In panels (a, c) the cavity is kept at the cold temperature ($\beta_{\text{C}}>\beta_{\text{A}}$), whereas in panels (b, d) it is the atom that has the colder temperature ($\beta_{\text{A}}>\beta_{\text{C}}$).
@@ III.3 Atom coupled to nonlinear cavity
Here, one of the subsystems is the atom, described by the Pauli matrix $\hat{\sigma}_{z}$ and characterized by the frequency $\omega_{\text{A}}$. The cavity is the other subsystem, and is described by the number operator $\hat{n}=\hat{a}^{\dagger}\hat{a}$, with $[\hat{a},\hat{a}^{\dagger}]=1$, and the cavity frequency $\omega_{\text{C}}$. We also include a Kerr nonlinearity, parametrized by $u_{\text{C}}$, which plays the role of effective interactions between cavity photons in this system. The atom and the cavity exchange photon quanta through the weak tunneling term $g(\hat{a}\hat{\sigma}_{+}+\hat{a}^{\dagger}\hat{\sigma}_{-})$ with $g\ll\hbar\omega_{\text{A,C}}$ and where $\hat{\sigma}_{+}$ and $\hat{\sigma}_{-}$ are the raising and lowering operators of the atom states. Note that, in the language of Eq. (20), we choose $\omega=0$ since any external driving frequency can be incorporated in the laser detuning—replacing $\omega_{\mathrm{C}}$—in the rotating frame [41]. For $u_{\text{C}}=0$, the Hamiltonian (30) reduces to the Jaynes-Cummings Hamiltonian [58].
@@
To relate to the theoretical framework of Sec. III.1, we now identify the tunneling operator $\hat{A}=g\hat{a}\hat{\sigma}_{+}$, which we use to calculate the zero-frequency amplitudes $\mathcal{C}_{\rightleftarrows}^{(i)}(0)$ in Eq. (28). Note that a possible observable $\hat{Q}$ with which a current and its noise may be defined according to (22) could here be the atomic occupation $\hat{Q}=\frac{\hat{\sigma}_{z}+1}{2}$, satisfying $[\hat{Q},\hat{A}]=\hat{A}$. The result for $\mathcal{C}_{\rightleftarrows}^{(i)}(0)$ is shown in Fig. 5, where we plot $\mathcal{C}_{\rightleftarrows}^{(i)}(0)$ for different values of the non-linearity $u_{\text{C}}$. Since we are free to choose the L subsystem, in panels (a, c) the atom is the subsystem considered at two different temperatures, whereas in panels (b, d) the cavity is. In general, Fig. 5 shows that the thermodynamic cost and the energetic cost are very similar to each other and that there is no hierarchy between them, namely it depends on the specific parameters whether the thermodynamic or the energetic cost is larger 333 This can for example be seen in panel (b), where, in the absence of Kerr non-linearity the thermodynamic cost is larger than the energetic one, but in the presence of such a non-linearity the opposite holds true.. We furthermore see that—unsurprisingly—both thermodynamic and energetic constraints are trivially saturated at equal temperature. However, for sizable temperature biases, considering the atom or the cavity at different temperatures affects the constraints.
@@
Panels (a, c) illustrate the case where the cavity is always taken at the cold temperature, while we compare the rates for the atom being taken at two different temperatures. We see that the rate $\Gamma_{\rightarrow}$ has a nontrivial constraint via the thermodynamic and energetic cost: for all temperature biases, the response contribution $\mathcal{C}^{(\mathrm{Resp})}_{\rightarrow}$ is positive and it thereby constrains the thermodynamic and energetic costs non-trivially. In contrast, for the rate $\Gamma_{\leftarrow}$, the response contribution is always negative, thereby not putting any nontrivial constraint on the energetic and the thermodynamic cost, which are positive by definition. This can be understood from the fact that the rate $\Gamma_{\leftarrow}$ characterizes the process in which the atom emits a photon into the cavity, requiring the atom to be in an excited state. However, if all allowed transitions start in states with energies that are more than a standard deviation away from the internal energy of the subsystem, here the atom, then the rate response is negative $\mathcal{W}^{\text{(Resp)}}\leq 0$, and the constraints become trivial, see Appendix B for a detailed proof.
In panels (b, d), we show results where the atom is always kept at the cold temperature, while the cavity can be taken at two different temperatures. Here, we see that the bounds are trivial in the absence of the non-linearity, $u_{\mathrm{C}}=0$. Instead, a nonvanishing cavity non-linearity allows one to approach both thermodynamic and energetic constraints also at large temperature biases for both rates $\Gamma_{\rightleftarrows}$. This can be understood from how the nonlinearity affects the constraints (29). There are two effects of the nonlinearity: On the one hand, finite values of $u_{\text{C}}$ break the degeneracy of the atom-cavity transitions, making the atom couple only to two consecutive cavity states. On the other hand, a large $u_{\text{C}}$ increases the energy spacing of the cavity. This feature reduces the number of states that have non negligible occupation at finite cavity temperature $\beta_{\text{C}}\neq 0$. These two aspects effectively reduce the size of the cavity (which here plays the role of the system on which the two-point measurement is performed, see Sec. II), moving it further away from the thermodynamic limit and thereby making the bounds more constraining.
III.4 Coupled fermionic rings
@@
Figure 6: Thermodynamic cost (dashed lines), energetic cost (dotted lines), and rate response (solid lines) for two (U and D) coupled fermionic rings with sizes $L_{\text{D}}=3$ and $L_{\text{U}}=10$. The costs are plotted vs the inverse temperature difference $\beta_{\text{D}}-\beta_{\text{U}}$. The ring hopping parameters, $t_{\text{D}}$ and $t_{\text{U}}$, and the driving frequency are fixed as $t_{\text{U}}=\frac{4}{5}t_{\text{D}}=2\hbar\omega$. The upper ring charging energy $u_{\text{U}}=0$ is also fixed, while the down ring charging energy is taken as $u_{\text{D}}/2t_{\text{D}}=0,10$ for curves in blue and red, respectively. In panels (a, c) the upper, larger ring is kept at the cold temperature ($\beta_{\text{U}}>\beta_{\text{D}}$), whereas in panels (b, d) it instead is the down, smaller ring that is colder ($\beta_{\text{L}}>\beta_{\text{U}}$).
@@
Figure 7: Thermodynamic cost (dashed lines), energetic cost (dotted lines), and rate response (solid lines) for the two coupled fermionic rings with charging energies $u_{\text{D}}/2t_{\text{D}}=10,u_{\text{U}}=0$, plotted vs the inverse temperature difference $\beta_{\text{U}}-\beta_{\text{D}}$. The lower ring has fixed size $L_{\mathrm{D}}=3$, and the upper ring has sizes $L_{\text{U}}=4,10$ for the curves in blue and red, respectively. All other parameters are the same as in Fig. 6. In panels (a, c), the upper, larger ring is kept at the cold temperature ($\beta_{\text{U}}>\beta_{\text{D}}$), whereas in panels (b, d) the lower, smaller ring is kept fixed at the cold temperature ($\beta_{\text{D}}>\beta_{\text{U}}$).
@@
with $g\ll t_{\alpha}$. Here, we identify the tunneling operator $\hat{A}=g\hat{c}^{\dagger}_{\text{D},0}\hat{c}_{\text{U},0}$, which transfers a fermion from the upper ring (U) to the lower ring (D). Then, using this tunneling operator $\hat{A}$, we calculate the amplitudes $\mathcal{C}_{\rightleftarrows}^{(i)}(\omega)$ of Eq. (28). Note that a possible observable $\hat{Q}$ could here be the number of fermions in one ring, e.g. $\hat{Q}=\hat{N}_{\text{D}}$, which satisfies $[\hat{Q},\hat{A}]=\hat{A}$. The result for $\mathcal{C}_{\rightleftarrows}^{(i)}(\omega)$ is shown in Fig. 6, where we plot $\mathcal{C}_{\rightleftarrows}^{(i)}(\omega)$ at different values of the charging energy $u_{\text{D}}$ in the lower ring, and in Fig. 7 for different sizes $L_{\text{U}}$ of the upper ring. Note that the frequency $\omega$ of the tunneling Hamiltonian can emerge from a potential bias between the rings after a gauge transformation [27, 60].
Similarly to the atom-cavity system considered in Sec. III.3, the charging energy $u_{\text{D}}$ influences the thermodynamic and energetic constraints in the system of coupled rings. In panels (a, c) of Fig. 6, we show results where the lower ring, namely the one displaying finite interaction, is the one that is considered at two different temperatures when comparing out-of-equilibrium rates. This produces nontrivial constraints for both $\Gamma_{\rightleftarrows}$, similar to the case of the atom-cavity system, where the cavity with nonlinearity $u_{\mathrm{C}}$ is taken at two different temperatures, as shown in Fig. 5 (b, d). In contrast, as shown in panels (b, d) of Fig. 6, for the case where the upper ring is taken at two different temperatures only, the rate $\Gamma_{\leftarrow}$ is non-trivially constrained, as long as the charging energy $u_{\text{U}}=0$ vanishes.
@@
Furthermore, as discussed in Sec. II.4, increasing the size of the ring that is considered at two different temperatures (namely corresponding to the one on which the two-point measurement is performed, see Sec. II), when comparing out-of-equilibrium rates, weakens the constraints, as is seen in panels (b, d) of Fig. 7. In this example, the tunneling rates scale inversely with the ring size, i.e. with a scaling function $f$ on the form $f(L_{\text{U}})\sim 1/L_{\text{U}}$ [see below Eq. (19)], whereas the extensive quantities entering the thermodynamic and energetic costs scale linearly with $L_{\text{U}}$. Thus, while the costs are essentially unaffected upon increasing $L_{\text{U}}$, the rate responses instead decrease. By contrast, if the ring that is considered at two different temperatures has a fixed size, both costs and rate responses scale as $1/L_{\text{U}}$ because of the scaling of the tunneling rates alone, see panels (a, c). This feature implies that, as long as the size of the subsystem considered at two different temperatures remains small, there are nontrivial constraints on the tunneling rates, irrespective of the size of the subsystem with fixed temperature.
IV Conclusions